OpenStudy (icalibear):

On January 1, 2010, Chessville has a population of 50,000 people. Chessville then enter a period of population growth. Its population increases 7% each year. On the same day, Checkersville has a population of 70,000 people. Checkersville starts to experience a population decline. Its population decreases 4% each year. During what year will the population of Chessville first exceed that of Checkersville? Please show your work and explain your steps in the simplest way possible. :)

OpenStudy (anonymous):

50,000(1.07)^t = 70,000(1-.04)^t = 70,000(.96)^t (1.07/.96)^t=(70,000/50,000)=1.4 t ln (1.07/.96) = ln(1.4) .10848 t = .33647 t=3.102 So population will be equal in 3.102 years

OpenStudy (anonymous):

Wow that was fast.

OpenStudy (icalibear):

Um can you explain what you just did please XD

OpenStudy (anonymous):

lol i googled it